OpenStudy (anonymous):

Two weather tracking stations are on the equator 146 miles apart. A weather balloon is located on a bearing of N 35°E from the western station and on a bearing of N 23°E from the eastern station. How far is the balloon from the western station? PLEASE HELP!!

OpenStudy (anonymous):



OpenStudy (anonymous):

that goes with it!

OpenStudy (anonymous):

 so all you need to do is find the interior angle 180 - 35 = 145 then find the other angle = 180- (145+23) = 12 Use sinelaw \[\frac{ 146 }{ \sin12 } = \frac{ x }{ \sin 113 }\]